OpenStudy (anonymous):

Simplify i^38 i -1 -i 1

OpenStudy (anonymous):

divide \(38\) by 4 and take the integer remainder i.e. the remainder is 2 that makes \(i^{38}=i^2=-1\)

OpenStudy (anonymous):

so B? thank youuu

OpenStudy (anonymous):

ok i don't get why you divided 38 by 4.. @satellite73

jimthompson5910 (jim_thompson5910):

You could do it the long way \[\Large i^{38} = i^{36+2}\] \[\Large i^{38} = i^{36}*i^{2}\] \[\Large i^{38} = i^{4*9}*i^{2}\] \[\Large i^{38} = (i^{4})^{9}*i^{2}\] \[\Large i^{38} = (1)^{9}*(-1)\] \[\Large i^{38} = 1*(-1)\] \[\Large i^{38} = -1\] But satellite73's shortcut is much simpler

OpenStudy (anonymous):

still dont get it :(

OpenStudy (anonymous):

im gonna post another question similar.

jimthompson5910 (jim_thompson5910):

\[\Large i^0 = 1\] \[\Large i^1 = i\] \[\Large i^2 = -1\] \[\Large i^3 = -i\] \[\Large i^4 = 1\] Notice how i^0 = i^4. So the pattern repeats every 4 times

jimthompson5910 (jim_thompson5910):

So that's why he divided by 4. The remainder tells you where in the pattern you are (no matter how huge the exponent gets)